# Angular momentum of vacuum bubbles in a first-order phase transition

### v1

@@ Abstract
The formation of primordial black holes (PBHs) during a first-order phase transition (FOPT) in a dark sector has been of recent interest. A quantity that characterizes a black hole is its spin. We carry out the first step towards determining the spin of such PBHs, by calculating the spin of spherical false vacuum bubbles induced by cosmological perturbations. The angular momentum is given by the product of density and velocity perturbations. We carefully track the evolution of background quantities and calculate the transfer functions during the FOPT. We find that the dimensionless spin parameter $s=J/(G_{\rm N}M^{2})$ of false vacuum bubbles of mass $M$ and angular momentum $J$, take a wide range of values from ${\cal{O}}(10^{-3})$ to ${\cal{O}}(10^{3})$ for FOPTs between 10 keV and 100 GeV and a dark sector that is 0.1 to 0.4 times cooler than the visible sector. We also find a scaling relation between the root-mean-square value of the spin, the FOPT time scale, the bubble wall velocity, and the dark sector-to-visible sector temperature ratio.
@@ 1 Introduction
A well-known result in general relativity (GR) is that black holes can be characterized by their mass and spin. As discussed extensively in the literature, the angular momentum of primordial black holes (PBHs) is strongly related to the black hole production mechanism. A common scenario is that enhancements in the primordial curvature power spectrum increase the probability of obtaining rare density peaks above the critical threshold for collapse [1, 2, 3, 4]. Computations of the spin of PBHs have been performed adopting various hypotheses, where they predict spin values in the range $s\simeq 10^{-3}\sim 10^{-2}$ [5, 6, 7, 8]. Alternatively, we focus on the scenario that PBHs are formed from a cosmological dark first-order phase transition (FOPT) without any enhancement in the primordial power spectrum.
@@
The PBH may be formed from the direct collapse of false vacuum (FV) bubbles [9]. Another mechanism involves the formation of intermediate compact objects called Fermi balls (FBs), which are composed of degenerate dark fermions [10, 11]. The FBs cool down by emitting light particles that may eventually trigger the collapse into PBHs. Both mechanisms efficiently produce subsolar mass FBs/PBHs for FOPT energy scales from $\mathcal{O}({\rm keV})$ to $\mathcal{O}(100\,{\rm GeV})$. In this work, the fundamental hypothesis in generating PBH spin is the idea that cosmological perturbations induce nonzero angular momentum. The cosmological perturbations are Gaussian random variables whose initial values are ultimately determined by the primordial curvature power spectrum. We show that the angular momentum and spin are Gaussian random variables and thus their distributions are determined by their root-mean-square (RMS) values. We restrict our calculation of the spin to the case of false vacuum (FV) bubbles for a number of reasons. In the case of direct collapse to PBHs, the process is nonlinear and may require techniques beyond perturbation theory. Similarly, in the case where FBs are formed, the emission of light species carry away angular momentum from the FV bubble. Tracking the subsequent evolution of the angular momentum is beyond the scope of our work.
@@
In order to realize the FOPT scenario, which is an essential ingredient in the novel PBH formation mechanism that we feature in this work, it is necessary to introduce dark sector dynamics, since phase transitions in the Standard Model are known to be smooth crossovers. In particular, we consider a model framework with a dark scalar $\phi$ that triggers the FOPT. For the case where the intermediate step of forming Fermi balls is present, we also require the presence of a dark Dirac fermion $\chi$, which will serve as the fundamental building block for Fermi balls. The size distribution of FV bubbles during the FOPT has been discussed in Ref. [12], from which the mass and abundance of FB/PBH can be determined. In our work, we take the size of the FV sphere to be the critical radius at the time of percolation, within which no true vacuum bubble can nucleate.
@@
We point out that the determination of the PBH spin, as well as its distribution, has been the subject of active investigation in the literature. Based on the seminal work [13] on the critical behavior of BH masses, arising from the evolution of a family of ingoing packets of a scalar field, Ref. [14] considered a radiation fluid characterized by the size of the density fluctuation and angular momentum. The main result of [14] is the presence of a power-law relation not only for the BH mass but also for the angular momentum. Building upon this work, Ref. [15] obtained the mass and angular momentum distribution by assuming that the initial density fluctuation is Gaussian, while the parameter describing the initial rotation was assumed to be flat. The resulting spin is around $s<0.4$; on the other hand, accretion effects on the mass and spin of the BHs were shown to be negligible. In the case where the PBH spin is induced by cosmological perturbations Ref. [5] calculated the RMS value of the PBH spin, assuming that the PBHs are formed from the collapse of rare spherically symmetric density peaks. The spin of the PBH was shown to be a product of density and velocity perturbations, making it a second-order quantity in perturbations. Assuming a narrow enhancement in the primordial curvature power spectrum, they obtained $s\simeq 0.01$, for PBHs comprising a sizable fraction of dark matter. In contrast, Ref. [6] assumed that the proto-object that will eventually collapse into PBHs is ellipsoidal, and the distribution of the ellipsoidal configurations is determined by peak theory [16]. They showed that the PBH spin is a first-order quantity in perturbations and obtained spin parameters at the percent level, given some assumptions in the enhancement in the curvature power spectrum. Ref. [7] followed the formalism laid down in Ref. [6], and provided some improvements in the analysis of the latter. Both approaches assumed that the initial PBH spin corresponds to the value at turn around time, i.e., the moment when the overdensity begins to collapse; in Ref. [17], they pointed out that the turn around time does not necessarily occur at the moment when the (enhanced) perturbation mode enters the horizon. As a result, Ref. [17] predicted an RMS spin value of $\simeq 4.0\times 10^{-3}(M/M_{H})^{-1/3}$, where $M_{H}$ is the mass within the Hubble horizon at horizon entry of the enhanced perturbation. In addition, Ref. [17] pointed out that the second-order approach in Ref. [5] may give rise to spin estimates that are comparable with the first-order approach. In Ref. [18], PBHs are formed from the collapse of domain walls and the angular momentum arises from the tidal forces from the surrounding radiation fluid acting on the walls, obtaining $s\simeq 10^{-4}(M/30M_{\odot})^{-1}$. Finally, recent work in Ref. [8] used the result from Ref. [6], and combined it with the curvature perturbation obtained in Ref. [19, 20], where they obtained a spin value of around $s\simeq\mathcal{O}(10^{-3})$. The curvature perturbation is generated from the difference in nucleation times of different points in space, thereby generating regions that contain more vacuum energy than others. We point out that the applicability of the result from Ref. [6] assumes that any modification in the curvature power spectrum should be made at the time when all perturbations are superhorizon.
@@
The rest of the paper will be devoted to the task of calculating the spin of FV bubbles at the percolation time, during a FOPT in the dark sector, induced by cosmological perturbations. In Section 2 we first introduce the dark sector model and the quartic scalar potential, before proceeding with the discussion of angular momentum in GR and in the specific case of a perturbed FRW spacetime in Section 3. Following Ref. [21], we present the evolution equations for cosmological perturbations in the uniform Hubble gauge in Section 4. We separately consider the period from the time when all modes are superhorizon up to the critical point, and the period of FOPT between the critical point and the percolation time. Determining the evolution of the background quantities and the perturbations during the FOPT are crucial to our task, and we provide a detailed discussion of this part of the calculation in Section 4.2. We present our main results, estimates for the FV spin, and benchmark points in Section 5. Finally, we conclude in Section 6.
@@ 2 Dark first-order phase transitions
In the case where PBHs are formed from the direct collapse of FV bubbles, it is sufficient to consider only a real scalar field $\phi$ in the dark sector. The scalar field triggers the FOPT by acquiring a nonzero expectation value at some critical point. On the other hand, if we consider the formation of Fermi balls, we also need to add Dirac dark fermions $\chi$ in the particle spectrum. These fermions serve as the fundamental component of Fermi balls. Then the dark sector effective Lagrangian density can be written as
@@
The form of the effective potential may arise from a more fundamental theory, possibly with more particle species, where the usual procedure of accounting for the zero-temperature and finite-temperature contributions to the effective potential at one-loop order, and taking the high-temperature limit are carried out, as in Refs. [22, 23]. The quantity $g_{\rho}(\phi)$ just counts the number of effective relativistic degrees of freedom in the dark sector for a given classical field configuration $\phi$, where bosons and fermions are respectively weighted by $1$ and $7/8$. Meanwhile, $\Delta b$ counts the number of degrees of freedom that are relativistic in the false vacuum but not in the true vacuum (TV). For simplicity, we take
@@
Note that the Yukawa term in Eq. (2.1) is mainly relevant for the Fermi ball scenario of PBH formation, but it can also be present in the scenario where PBHs are directly formed from FV bubbles. Firstly, the Yukawa term allows the filtered DM scenario to be realized, where the $\chi$ species are trapped inside FV bubbles. A necessary condition for this to occur is that the $\chi$ mass difference $\Delta m_{\chi}=g_{\chi}\phi_{+}$ between the TV and FV is much larger than the typical kinetic energy of $\chi$, roughly given by $T$, so that the flux of $\chi$ particles through the bubble walls is exponentially suppressed. Secondly, the Yukawa term provides an attractive force between the $\chi$ particles to trigger an instability that leads to the formation of PBH from FBs. Finally, under certain conditions, the Yukawa interaction ensures that the $\chi$ and $\phi$ species are tightly coupled with each other. The necessary condition for this to occur can be derived by requiring that interactions involving $\chi$ and $\phi$, e.g. $\chi\bar{\chi}\rightarrow\phi\phi$, occur at a much faster rate than the Hubble expansion, so that we have
@@
The dynamics of the FOPT is determined by tracking the cosmological evolution of the plasma, in the presence of both SM and dark sector degrees of freedom, treating each component as a fluid with a known equation of state. For the dark sector we can calculate the thermodynamic quantities—free energy, pressure, entropy, and energy—in the false and true vacuum using the effective potential Eq. (2.2). These are then given respectively by (cf. [24, 25, 26, 27])
@@
To simplify our analysis, we match the thermodynamic quantities with the bag model [25], by expanding $\Delta V(T)$ about the critical temperature $T_{\rm c}$, up to linear order. By definition, the critical temperature is the moment at which $\Delta p=0$, so that $\Delta V(T_{\rm c})=0$. In terms of the parameters in the quartic potential,
@@
Meanwhile, the derivative of $\Delta p$ is discontinuous at $T=T_{\rm c}$. This is simply attributed to the presence of a latent heat, which is a defining characteristic of a first-order phase transition. Denoting the latent heat at the critical temperature by $L_{\rm c}$, where
@@
$\Delta p(T)\approx-L_{\rm c}\left(1-\frac{T}{T_{\rm c}}\right)\,,\quad\Delta s(T)\approx\frac{L_{\rm c}}{T_{\rm c}}\,,\quad\Delta\rho(T)\approx L_{\rm c}\,.$ (2.15)
In terms of the parameters in the effective potential, we have
@@
Since $0\leq T_{0}^{2}/T_{\rm c}^{2}\leq 1$, there must be a lower bound for the latent heat, i.e.
@@ 2.1 Physical conditions
We further impose physical conditions on the thermodynamic quantities, to restrict the class of effective potentials that will be the subject of this study. These conditions should apply for all temperatures within $T_{0}\leq T\leq T_{\rm c}$. In the bag model, $\Delta p=-L_{\rm c}(1-T/T_{\rm c})\leq 0$ is automatically satisfied, which ensures that the TV regions exert a net outward pressure on the surrounding FV region, to allow nucleated TV bubbles to expand. Furthermore, we require $\Delta\rho\geq 0$ since the FV has a higher energy than the TV. This translates to
$\frac{\pi^{2}}{30}g_{\rho}T_{0}^{4}\geq L_{\rm c}\,,$ (2.19)
so that
$\frac{L_{\rm c}}{T_{\rm c}^{4}}\leq\frac{\pi^{2}}{30}g_{\rho}T_{0}^{4}/T_{\rm c}^{4}.$ (2.20)
It can be easily seen that the rescaled latent heat $L_{\rm c}/T_{\rm c}^{4}$ cannot be increased indefinitely. Suppose we take a large value of $L_{\rm c}/T_{\rm c}^{4}$, so that $T_{0}^{2}/T_{\rm c}^{2}\rightarrow 1$ according to Eq. (2.17). But the right hand side of Eq. (2.20) suggests that $L_{\rm c}/T_{\rm c}^{4}$ cannot exceed $g_{\rho}\pi^{2}/30$. Analogously, the greatest lower bound for the rescaled latent heat must be larger than the one prescribed in Eq. (2.18) which leads to $T_{0}^{2}/T_{\rm c}^{2}=0$, because of Eq. (2.20). In addition, the square of the sound speed in the TV, taken in isolation, differs from that of a radiation fluid, and is given by
$\boxed{c_{s\rm,TV}^{2}=\frac{dp_{\rm TV}/dT}{d\rho_{\rm TV}/dT}\simeq\frac{1}{3}\left[1-\frac{15}{\pi^{2}g_{\rho}}\left(\frac{L_{\rm c}}{T_{\rm c}^{4}}\right)\left(\frac{T}{T_{\rm c}}\right)^{-3}\right]}\,.$ (2.21)
One can check that the physical condition Eq. (2.20) guarantees that $c_{s\rm,TV}^{2}$ remains positive and subluminal. In general, Eq. (2.20) can be recast using Eq. (2.17) as
$\gamma u^{3}-u^{2}+\gamma bu+a<0,\quad\gamma\equiv\left(\frac{30}{\pi^{2}g_{\rho}}\right)^{1/2},\quad u\equiv\left(\frac{L_{\rm c}}{T_{\rm c}^{4}}\right)^{1/2},$ (2.22)
$a\equiv\frac{8x_{C}(A+x_{C})^{3}}{\lambda^{3}},\quad b\equiv\frac{8A(A+x_{C})^{3}}{\lambda^{3}}.$ (2.23)
For a given pair of $A$ and $x_{C}$, the problem is reduced to solving the roots of the cubic polynomial. We want the case where we have two positive real roots, which will determine the allowed range of $L_{\rm c}/T_{\rm c}^{4}$. It is easy to show that there is always one negative real root, while there can be two or no positive real roots.
We find it convenient to consider
@@
as fundamental parameters. Once these parameters are specified, the effective potential is known. In choosing our benchmark points, we take $A=10^{-1},10^{-2},10^{-3},10^{-4}$. As for the other parameters, we turn to the physical conditions for guidance. It turns out that there is an absolute minimum value of $\lambda=\lambda_{0}$, below which there exists no physical parameter region in the $(L_{\rm c}/T_{\rm c}^{4},x_{C})$ plane. This can be obtained by setting the discriminant of the cubic polynomial in Eq. (2.22) to zero, i.e.
$\frac{1}{\lambda^{3}}$$\cdot$ $\left[c_{2}\left(\frac{1}{\lambda^{3}}\right)^{2}+c_{1}\left(\frac{1}{\lambda^{3}}\right)+c_{0}\right]=0,$ (2.24)
$c_{2}$$\equiv$ $55296A^{3}(A+x_{C})^{9}\gamma^{6},$ (2.25)
$c_{1}$$\equiv$ $-1728(A+x_{C})^{6}\left(A^{2}-18Ax_{C}-27x_{C}^{2}\right)\gamma^{4},$ (2.26)
$c_{0}$$\equiv$ $-864x_{C}(A+x_{C})^{3}\gamma^{2},$ (2.27)
and then setting $x_{C}=0$. We find that
$\lambda_{0}=\left(32A^{4}\gamma^{2}\right)^{1/3}\simeq 0.129\left(\frac{A}{0.1}\right)^{4/3}\left(\frac{4.5}{g_{\rho}}\right)^{1/3}.$ (2.28)
For given values of $A$ and $\lambda$, the values of $x_{C}$ and $L_{\rm c}/T_{\rm c}^{4}$ are chosen to lie in the physical region, such that Eq. (2.22) is satisfied. On the other hand, if we choose $A$ and $\lambda$, then Eq. (2.24) becomes a polynomial equation for the maximum $x_{C}$ that can accommodate physical values of $L_{\rm c}/T_{\rm c}^{4}$.
@@
Figure 1: The physical region for $A=0.1,0.001$, and two values of $\lambda/\lambda_{0}=1.42,2.34$. The green curves correspond to contours of constant $T_{0}/T_{\rm c}$. For all cases, $g_{\rho}=4.5$.
As an illustration, in Fig. 1 we show the physical region as unshaded, for $A=0.1,0.001$, choosing two values of $\lambda/\lambda_{0}$ for each: 1.42 and 2.34. The black line corresponds to $T_{0}=0$, so that the region above it leads to $T_{0}^{2}<0$, and is thus forbidden. The green contours correspond to contours of constant $T_{0}/T_{\rm c}=0.1,0.2,...,$ to $0.9$, indicated by solid curves. We also show contours where $T_{0}/T_{\rm c}$ is very close to unity, with $T_{0}/T_{\rm c}=1-10^{-2},1-10^{-3},...$, indicated by dashed curves. Meanwhile, the shaded red region corresponds to FOPT scenarios that are excluded by the physical condition $\rho_{\rm TV}<0$, for all $T_{0}\leq T\leq T_{\rm c}$. The horizontal dashed line marks the maximum $x_{C}$ value allowed under this physical condition, for fixed $A$ and $\lambda$. We also observe that a smaller $x_{C}$ leads to a wider range of allowed $L_{\rm c}/T_{\rm c}^{4}$ values. The maximum range of physical $L_{\rm c}/T_{\rm c}^{4}$ values is achieved when $x_{C}=0$, for fixed $A$ and $\lambda$.
Note that a larger $\lambda$ leads to a larger area occupied by the physical region, as well as an increase in the maximum $x_{C}$. Smaller values of $L_{\rm c}/T_{\rm c}^{4}$ are accommodated, opening up a wider range of physical rescaled latent heat values.

### v2

@@ Abstract
The formation of primordial black holes (PBHs) during a first-order phase transition (FOPT) in a dark sector has been of recent interest. A quantity that characterizes a black hole is its spin. We carry out the first step towards determining the spin of such PBHs, by calculating the spin of spherical false vacuum bubbles induced by cosmological perturbations. The angular momentum is given by the product of density and velocity perturbations. We carefully track the evolution of background quantities and calculate the transfer functions during the FOPT. We find that the dimensionless spin parameter $s=J/(G_{\rm N}M^{2})$ of false vacuum bubbles of mass $M$ and angular momentum $J$, take a wide range of values from ${\cal{O}}(10^{-5})$ to ${\cal{O}}(10)$ for FOPTs between 10 keV and 100 GeV and a dark sector that is 0.1 to 0.4 times cooler than the visible sector. We also find a scaling relation between the root-mean-square value of the spin, the FOPT time scale, the bubble wall velocity, and the dark sector-to-visible sector temperature ratio.
@@ 1 Introduction
A well-known result in general relativity (GR) is that black holes can be characterized by their mass and spin. As discussed extensively in the literature, the angular momentum of primordial black holes (PBHs) is strongly related to the black hole production mechanism. A common scenario is that enhancements in the primordial curvature power spectrum increase the probability of obtaining rare density peaks above the critical threshold for collapse [Press:1973iz, Green:2004wb, Young:2020xmk, Yoo:2020dkz]. Computations of the spin of PBHs have been performed adopting various hypotheses, where they predict spin values in the range $s\simeq 10^{-3}\sim 10^{-2}$ [Mirbabayi:2019uph, DeLuca:2019buf, Harada:2020pzb, Banerjee:2023qya]. Alternatively, we focus on the scenario that PBHs are formed from a cosmological dark first-order phase transition (FOPT) without any enhancement in the primordial power spectrum.
@@
The PBH may be formed from the direct collapse of false vacuum (FV) bubbles [Baker:2021nyl]. Another mechanism involves the formation of intermediate compact objects called Fermi balls (FBs), which are composed of degenerate dark fermions [Gross:2021qgx, Kawana:2021tde]. The FBs cool down by emitting light particles that may eventually trigger the collapse into PBHs. Both mechanisms efficiently produce subsolar mass FBs/PBHs for FOPT energy scales from $\mathcal{O}({\rm keV})$ to $\mathcal{O}(100\,{\rm GeV})$. In this work, the fundamental hypothesis in generating PBH spin is the idea that cosmological perturbations induce nonzero angular momentum. The cosmological perturbations are Gaussian random variables whose initial values are ultimately determined by the primordial curvature power spectrum. We show that the angular momentum and spin are Gaussian random variables and thus their distributions are determined by their root-mean-square (RMS) values. We restrict our calculation of the spin to the case of false vacuum (FV) bubbles for a number of reasons. In the case of direct collapse to PBHs, the process is nonlinear and may require techniques beyond perturbation theory. Similarly, in the case where FBs are formed, the emission of light species carry away angular momentum from the FV bubble. Tracking the subsequent evolution of the angular momentum is beyond the scope of our work.
@@
To realize the FOPT scenario, which is an essential ingredient in the novel PBH formation mechanism that we feature in this work, it is necessary to introduce dark sector dynamics, since phase transitions in the Standard Model are known to be smooth crossovers. In particular, we consider a model framework with a dark scalar $\phi$ that triggers the FOPT. For the case in which Fermi balls form in an intermediate step, we also require dark Dirac fermion $\chi$, which serve as the fundamental building block for Fermi balls. The size distribution of FV bubbles during the FOPT has been discussed in Ref. [Lu:2022paj], from which the mass and abundance of FB/PBH can be determined. In our work, we take the size of the FV sphere to be the critical radius at the time of percolation, within which no true vacuum bubble can nucleate.
@@
We point out that the determination of the PBH spin, as well as its distribution, has been the subject of active investigation in the literature. Based on the seminal work [Choptuik:1992jv] on the critical behavior of BH masses, arising from the evolution of a family of ingoing packets of a scalar field, Ref. [Baumgarte:2016xjw] considered a radiation fluid characterized by the size of the density fluctuation and angular momentum. The main result of [Baumgarte:2016xjw] is the presence of a power-law relation not only for the BH mass but also for the angular momentum. Building upon this work, Ref. [Chiba:2017rvs] obtained the mass and angular momentum distribution by assuming that the initial density fluctuation is Gaussian, while the parameter describing the initial rotation was assumed to be flat. The resulting spin $s\lesssim 0.4$; on the other hand, accretion effects on the mass and spin of the BHs were shown to be negligible. In the case where the PBH spin is induced by cosmological perturbations Ref. [Mirbabayi:2019uph] calculated the RMS value of the PBH spin, assuming that the PBHs are formed from the collapse of rare spherically symmetric density peaks. The spin of the PBH was shown to be a product of density and velocity perturbations, making it a second-order quantity in perturbations. Assuming a narrow enhancement in the primordial curvature power spectrum, they obtained $s\simeq 0.01$, for PBHs comprising a sizable fraction of dark matter. In contrast, Ref. [DeLuca:2019buf] assumed that the proto-object that will eventually collapse into PBHs is ellipsoidal, and the distribution of the ellipsoidal configurations is determined by peak theory [heavens1988tidal]. They showed that the PBH spin is a first-order quantity in perturbations and obtained spin parameters at the percent level, given some assumptions in the enhancement in the curvature power spectrum. Ref. [Harada:2020pzb] followed the formalism laid down in Ref. [DeLuca:2019buf], and provided some improvements in the analysis of the latter. Both approaches assumed that the initial PBH spin corresponds to the value at turn around time, i.e., the moment when the overdensity begins to collapse; in Ref. [Harada:2013epa], they pointed out that the turn around time does not necessarily occur at the moment when the (enhanced) perturbation mode enters the horizon. As a result, Ref. [Harada:2013epa] predicted an RMS spin value of $\simeq 4.0\times 10^{-3}(M/M_{H})^{-1/3}$, where $M_{H}$ is the mass within the Hubble horizon at horizon entry of the enhanced perturbation. In addition, Ref. [Harada:2013epa] pointed out that the second-order approach in Ref. [Mirbabayi:2019uph] may give rise to spin estimates that are comparable with the first-order approach. In Ref. [Eroshenko:2021sez], PBHs are formed from the collapse of domain walls and the angular momentum arises from the tidal forces from the surrounding radiation fluid acting on the walls, obtaining $s\simeq 10^{-4}(M/30M_{\odot})^{-1}$. Finally, recent work in Ref. [Banerjee:2023qya] used the result from Ref. [DeLuca:2019buf], and combined it with the curvature perturbation obtained in Ref. [Liu:2022lvz, Elor:2023xbz], where they obtained a spin value of around $s\simeq\mathcal{O}(10^{-3})$. The curvature perturbation is generated from the difference in nucleation times of different points in space, thereby generating regions that contain more vacuum energy than others. We point out that the applicability of the result from Ref. [DeLuca:2019buf] assumes that any modification in the curvature power spectrum should be made at the time when all perturbations are superhorizon.
@@
The rest of the paper is devoted to the task of calculating the spin of FV bubbles at the percolation time, during an FOPT in the dark sector, induced by cosmological perturbations. In Section 2 we first introduce the dark sector model and the quartic scalar potential, before proceeding with the discussion of angular momentum in GR and in the specific case of a perturbed Friedmann–Robertson–Walker (FRW) spacetime in Section 3. Following Ref. [Schmid:1998mx], we present the evolution equations for cosmological perturbations in the uniform Hubble gauge in Section 4. We separately consider the period from the time when all modes are superhorizon up to the critical point, and the period of FOPT between the critical point and the percolation time. Determining the evolution of the background quantities and the perturbations during the FOPT are crucial to our task, and we provide a detailed discussion of this part of the calculation in Section 4.2. We present our main results, estimates for the FV spin, and benchmark points in Section 5. Finally, we conclude in Section 6.
@@ 2 Dark first-order phase transitions
In the case where PBHs are formed from the direct collapse of FV bubbles, it is sufficient to consider only a real scalar field $\phi$ in the dark sector. The scalar field triggers the FOPT by acquiring a nonzero expectation value at some critical point. On the other hand, to consider the formation of Fermi balls we include almost massless dark Dirac fermions $\chi$ in the particle spectrum. These fermions serve as the fundamental component of Fermi balls. Then the dark sector effective Lagrangian density can be written as
@@
The form of the effective potential may arise from a more fundamental theory, possibly with more particle species, where the usual procedure of accounting for the zero-temperature and finite-temperature contributions to the effective potential at one-loop order, and taking the high-temperature limit are carried out, as in Refs. [Dolan:1973qd, Anderson:1991zb]. The quantity $g_{\rho}(\phi)$ counts the number of effective relativistic degrees of freedom in the dark sector for a given classical field configuration $\phi$, where bosons and fermions are respectively weighted by $1$ and $7/8$. $\Delta b$ counts the number of degrees of freedom that are relativistic in the false vacuum but not in the true vacuum (TV). For simplicity, we take
@@
[…]
@@
The dynamics of the FOPT is determined by tracking the cosmological evolution of the plasma, in the presence of both SM and dark sector degrees of freedom, treating each component as a fluid with a known equation of state. For the dark sector we can calculate the thermodynamic quantities—free energy, pressure, entropy, and energy—in the false and true vacuum using the effective potential Eq. (2.2). These are then given respectively by (cf. [Leitao:2014pda, Espinosa:2010hh, Giese:2020rtr, Tenkanen:2022tly])
@@
By definition, the critical temperature is the moment at which $\Delta p=0$, so that $\Delta V(T_{\rm c})=0$. In terms of the parameters in the quartic potential,
@@
The derivative of $\Delta p$ is discontinuous at $T=T_{\rm c}$ and is simply attributed to the presence of latent heat, which is a defining characteristic of an FOPT. Denoting the latent heat at the critical temperature by $L_{\rm c}$, where
@@
Since $0\leq T_{0}^{2}/T_{\rm c}^{2}\leq 1$, there must be a lower bound for the latent heat, i.e.,
@@ 2.1 Physical conditions
We further impose physical conditions on the thermodynamic quantities, to restrict the class of effective potentials that will be the subject of this study. These conditions should apply for all temperatures within $T_{0}\leq T\leq T_{\rm c}$. Generally, $\Delta p=-\Delta V_{\rm eff}\leq 0$ is automatically satisfied, which ensures that the TV regions exert a net outward pressure on the surrounding FV region, to allow nucleated TV bubbles to expand. Furthermore, we require $\Delta\rho\geq 0$ since the FV has a higher energy than the TV. The energy difference between the two phases is a monotonically increasing function of $T$. On the other hand, we also need to ensure that $\rho_{\rm TV}(T)=g_{\rho}\pi^{2}/30~T^{4}-\Delta\rho(T)\geq 0$ for all $T_{0}\leq T\leq T_{\rm c}$. We also require that the square of the sound speed in the TV, taken in isolation,
$\boxed{c_{s\rm,TV}^{2}=\frac{dp_{\rm TV}/dT}{d\rho_{\rm TV}/dT}}\,$, (2.18)
be positive and subluminal. If $s_{\rm TV}=dp_{\rm TV}/dT\geq 0$, then a necessary condition for $c_{s\rm,TV}^{2}>0$ is $d\rho_{\rm TV}/dT>0$; this latter condition and $\rho_{\rm TV}(T_{0})\geq 0$ guarantee $\rho_{\rm TV}(T)\geq 0$. We find it convenient to consider
@@
as fundamental parameters. Once these parameters are specified, the effective potential is known. In choosing our benchmark points, we take $A=10^{-1},10^{-2},10^{-3},10^{-4}$. As for the other parameters, we turn to the physical conditions for guidance. If $\Delta\rho^{\prime}(T)>0$, it follows that if $\rho_{\rm TV}(T_{0})=0$,
$\frac{\pi^{2}}{30}g_{\rho}\left(\frac{T_{0}}{T_{\rm c}}\right)^{4}=\frac{\Delta\rho(T_{0})}{T_{\rm c}^{4}}\leq\frac{L_{\rm c}}{T_{\rm c}^{4}}.\,$ (2.19)
At $x_{C}=0$ we find that
$\frac{\pi^{2}}{30}g_{\rho}\left(\frac{T_{0}}{T_{\rm c}}\right)^{2}\left[1-\left(\frac{T_{0}}{T_{\rm c}}\right)^{2}\right]\leq\frac{8A^{4}}{\lambda^{3}}.\,$ (2.20)
In the range $0\leq T_{0}/T_{\rm c}\leq 1$, the maximum value of the left hand side of the inequality is $g_{\rho}\pi^{2}/120$. Then we define $\lambda_{0}$ as a reference scale for $\lambda$ via
$\frac{\pi^{2}}{120}g_{\rho}\equiv\frac{8A^{4}}{\lambda_{0}^{3}}\rightarrow\lambda_{0}\simeq 0.129\left(\frac{A}{0.1}\right)^{4/3}\left(\frac{4.5}{g_{\rho}}\right)^{1/3}\,.$ (2.21)
@@
Figure 1: The physical region is shaded in blue for $A=10^{-3},10^{-4}$, and two values of $\lambda/\lambda_{0}\simeq 1.33,1.72$. The green curves correspond to contours of constant $T_{0}/T_{\rm c}$, from 0.1 to 0.9, in steps of 0.1. The dashed green contour corresponds to $T_{0}/T_{\rm c}=0.99$. For all cases, $g_{\rho}=4.5$.
As an illustration, in Fig. 1 we show the physical region shaded in blue, for $A=10^{-3},10^{-4}$, choosing two values of $\lambda/\lambda_{0}$ for each: 1.32529 and 1.72359. The black line corresponds to $T_{0}=0$, so that the region above it leads to $T_{0}^{2}<0$, and is thus forbidden. The green contours correspond to contours of constant $T_{0}/T_{\rm c}=0.1,0.2,$..., to $0.9$, indicated by solid curves. The lone dashed curve in each panel corresponds to $T_{0}/T_{\rm c}=0.99$. We observe that a smaller $x_{C}$ leads to a wider range of allowed $L_{\rm c}/T_{\rm c}^{4}$ values. The maximum range of physical $L_{\rm c}/T_{\rm c}^{4}$ values is achieved when $x_{C}=0$, for fixed $A$ and $\lambda$. Note that a larger $\lambda$ leads to a larger area occupied by the physical region, as well as an increase in the maximum $x_{C}$. Smaller values of $L_{\rm c}/T_{\rm c}^{4}$ are accommodated, opening up a wider range of physical rescaled latent heat values.